OpenStudy (anonymous):

Give the slope of each line and determine whether the lines are parallel, perpendicular or neither.? 3x+5y=5 9x+15y=0

OpenStudy (anonymous):

convert them into the form y=mx+c where m is slope and c is the constant

OpenStudy (anonymous):

and for perpendicularity check m1*m2=-1 where m1= slope of 1st line m2=slope of second line

OpenStudy (anonymous):

m=-3/5; b=0, neither